# VSF: Simple, Efficient, and Effective Negative Guidance in Few-Step Image Generation Models By Value Sign Flip

### v5

@@ Acknowledgements
This work was supported by the NFRF under grant GR024801 and the CFI under grant GR024473. We also acknowledge Weathon Software (https://weasoft.com) for providing computing credits via Google Colab, and Lambda, Inc. (https://lambda.ai/) for computing credits via Lambda Cloud and Lambda Inference.
@@ Appendix D Details About the Metrics
To evaluate the scores of the generated images, we used LLaMA 4 Maverick, which has a very high image reasoning MMMU Yue et al. (2024) score, higher than Gemma 3 and even GPT-4o. We avoided using the same model (o3) for both evaluation and generation for cost control and to avoid bias within a model. We did not evaluate the quality of the generated images using models like ImageReward (Xu et al., 2023) or HPSv2 Wu et al. (2023) as in NASA or NAG, as current quality or human preference assessment models do not account for negative prompts; traditional methods usually aim for real-world generations Ye et al. (2025). Removing a key element from the positive prompt (e.g., removing the roof from a house) is likely to reduce perceived quality, since the result deviates from what is considered “normal,” even though that is the intended outcome. An example is shown in Figure 8 where removing the keys from the piano results in a much lower quality score, even though other parts of the image look the same. Both ImageReward and HPSv2 are built on top of image-text alignment models (CLIP (Radford et al., 2021) or BLIP (Li et al., 2022)), which will likely lead to a decreased score when the main object is missing a critical part. Thus, we also let the MLLM rate the image quality from 0-1 for each image and told it to ignore the abnormality of following the negative prompt. We did observe that MLLM can make mistakes, especially when there is ambiguity or when the unwanted element is hard to see. However, we believe that in general, under 400 images, the mistakes are minor. To compensate for this, we also did a human evaluation and a more negotiation-aware MLLM evaluation to cross-validate the LLaMA evaluation. Due to model provider stability issues, we used different MLLM providers for different portions of the experiment for the same model under the same config; this could have some limited impact on the stability of metrics.
@@ Appendix E Hyper-parameter Tunning
Although NAG (Chen et al., 2025a) also targeted negative concept avoidance, its primary focus was on its effects on improving generation quality (using words like “blurry” or “low quality” as a negative prompt). We believe the hyperparameters reported in their work were tuned with an emphasis on quality rather than negation handling. Therefore, we re-tuned their hyperparameters moderately and manually on guidance scale ($\phi$), blending factor ($\alpha$), and normalization factor ($\tau$). We will report experimental results on both original NAG (noted as NAG) and the improved hyperparameter version (noted as NAG Strong). The final hyperparameters used are $\phi=11,\alpha=0.5,\tau=5$. This pushes the NAG to the edge of acceptable visual quality.
@@ Appendix I External Baselines
In this paper, we focused on removing a critical component in the image. To further validate our negative guidance method in other areas, we also tested it on style avoidance. In Figure 5 (in main text), we show four examples, each of which is generated using the same seed. We can see that when prompted with famous artwork (e.g.,”A painting of Starry Night from the 1890s” or ”Mona Lisa oil painting”) but with a negative prompt of the artist’s name style, the generated image avoided the specific art style but kept the semantic meaning of the positive prompt. When prompted to give an old photo but not monochrome, the generated image is more like an old-style color photo, follows both non-monochrome and also not very bright (as old photos, even in color, are less vibrant). We find these examples interesting and think they can be used for machine unlearning, using a similar method as in (Gandikota et al., 2023).
@@
The first test used prompts containing the same object in both positive and negative form, with the goal of producing abstract art. This works by semi-canceling the main object, making it appear in an abstract form. Abstract styles are often disadvantaged in alignment settings, since reward models typically favor realistic or figurative outputs. VisionReward (Xu et al., 2025) encodes this bias through its scoring metric, and LAPIS (Maerten et al., 2025) reports that abstract paintings generally receive lower preference scores. Figure 15 shows that an abstract image gets a much lower score compared with a figurative one. As shown in the first two rows of Figure 18, the apple, people, and cat appear in abstract form, demonstrating a clear shift away from the default figurative tendency of the aligned models when VSF is applied. For the last image of a dog, we used “cute” as a negative prompt, which usually describes realistic objects, and we achieved a very abstract and artistic image. Figure 16 shows how abstract the image gets as the scale increases. More examples are shown in Figure 17.
@@
The model was finetuned using prompts from the dataset of all 3 models. We trained the model using QLoRA Dettmers et al. (2023) with rank of 8 and $r=8$, applied to query, key, value projections in both the vision encoder and language model with dropout of 0.1. Model is trained using $lr=5\times 10^{-5}$ (with warm up and decay), WeightDecay=0.1, BatchSize=16, Epoch=5. The dataset is split into train-val with a 90-10 ratio based on the prompt level splitting and aiming for balanced scores in each split. We treat 0.5 as False and calculate negative scores as a binary metric. Results are presented in Table 7 with comparison with the same model without finetuning and LLaMA-Maverick.
@@
Appendix S Use of LLMs in The Paper
In this work, we used LLM for paper-related work collection and consumption. It is also used to polish the paper language or provide feedback for writing/figures. It is also used to brainstorm before and during the project.

### v6

@@ Acknowledgements
This work was supported by the NFRF under grant GR024801 and the CFI under grant GR024473. The authors acknowledge Weathon Software (https://weasoft.com) and Lambda, Inc. (https://lambda.ai/) for providing computing resources via Google Colab and Lambda Cloud, respectively.
Following the completion of this research and paper submission, we became aware of NegPiP (https://github.com/hako-mikan/sd-webui-negpip/), a community project that utilizes an approach similar to ours. While this work was developed entirely independently from NegPiP, we recognize its contribution and suggest it as a relevant resource for readers.
@@ Appendix A Negation in Vision Language Models
Figure 6: Attention maps and intermediate images during the diffusion process. The leftmost column shows the final generated image (top) and an image generated without applying VSF scaling ($\alpha=0$, bottom). The top row on the right side displays the unnormalized attention values between image tokens and negative prompt tokens, while the bottom row shows the corresponding intermediate images at each timestep. The negative prompt is “umbrella.”
@@ Appendix D Details About the Metrics
To evaluate the scores of the generated images, we used LLaMA 4 Maverick, which has a very high image reasoning MMMU (Yue et al., 2024) score, higher than Gemma 3 and even GPT-4o. We avoided using the same model (o3) for both evaluation and generation for cost control and to avoid bias within a model. We did not evaluate the quality of the generated images using models like ImageReward (Xu et al., 2023) or HPSv2 (Wu et al., 2023) as in NASA or NAG, as current quality or human preference assessment models do not account for negative prompts; traditional methods usually aim for real-world generations (Ye et al., 2025). Removing a key element from the positive prompt (e.g., removing the roof from a house) is likely to reduce perceived quality, since the result deviates from what is considered “normal,” even though that is the intended outcome. An example is shown in Figure 8 where removing the keys from the piano results in a much lower quality score, even though other parts of the image look the same. Both ImageReward and HPSv2 are built on top of image-text alignment models (CLIP (Radford et al., 2021) or BLIP (Li et al., 2022)), which will likely lead to a decreased score when the main object is missing a critical part. Thus, we also let the MLLM rate the image quality from 0-1 for each image and told it to ignore the abnormality of following the negative prompt. We did observe that MLLM can make mistakes, especially when there is ambiguity or when the unwanted element is hard to see. However, we believe that in general, under 400 images, the mistakes are minor. To compensate for this, we also did a human evaluation and a more negotiation-aware MLLM evaluation to cross-validate the LLaMA evaluation. Due to model provider stability issues, we used different MLLM providers for different portions of the experiment for the same model under the same config; this could have some limited impact on the stability of metrics.
@@ Appendix E Hyper-parameter Tunning
Although NAG (Chen et al., 2025a) also targeted negative concept avoidance, its primary focus was on its effects on improving generation quality (using words like “blurry” or “low quality” as a negative prompt). We believe the hyperparameters reported in their work were tuned with an emphasis on quality rather than negation handling. Therefore, we re-tuned their hyperparameters moderately and manually on guidance scale ($\phi$), blending factor ($\alpha$), and normalization factor ($\tau$). We will report experimental results on both original NAG (noted as NAG) and the improved hyperparameter version (noted as NAG Strong). The final hyperparameters used are $\phi=11,\alpha=0.5,\tau=5$ for NAG Strong and $\phi=4,\alpha=0.125,\tau=2.5$ for original NAG. This pushes the NAG to the edge of acceptable visual quality.
@@ Appendix I External Baselines
In this paper, we focused on removing a critical component in the image. To further validate our negative guidance method in other areas, we also tested it on style avoidance. In Figure 4 (in main text), we show four examples, each of which is generated using the same seed. We can see that when prompted with famous artwork (e.g.,”A painting of Starry Night from the 1890s” or ”Mona Lisa oil painting”) but with a negative prompt of the artist’s name style, the generated image avoided any elements related to the style (including the town in the Starry Night) but kept the semantic meaning of the positive prompt. When prompted to give an old photo but not monochrome, the generated image is more like an old-style color photo, follows both non-monochrome and also not very bright (as old photos, even in color, are less vibrant). We find these examples interesting and think they can be used for machine unlearning, using a similar method as in (Gandikota et al., 2023).
@@
The first test used prompts containing the same object in both positive and negative form, with the goal of producing abstract art. This works by semi-canceling the main object, making it appear in an abstract form. Abstract styles are often disadvantaged in alignment settings, since reward models typically favor realistic or figurative outputs. VisionReward (Xu et al., 2025) encodes this bias through its scoring metric, and LAPIS (Maerten et al., 2025) reports that abstract paintings generally receive lower preference scores. Figure 15 shows that an abstract image gets a much lower score compared with a figurative one. As shown in the first two rows of Figure 17, the apple, people, and cat appear in abstract form, demonstrating a clear shift away from the default figurative tendency of the aligned models when VSF is applied. For the last image of a dog, we used “cute” as a negative prompt, which usually describes realistic objects, and we achieved a very abstract and artistic image. Figure 16 shows how abstract the image gets as the scale increases. More examples are shown in Figure LABEL:fig:abstract2.
@@
The model was finetuned using prompts from the dataset of all 3 models. We trained the model using QLoRA (Dettmers et al., 2023) with rank of 8 and $r=8$, applied to query, key, value projections in both the vision encoder and language model with dropout of 0.1. Model is trained using $lr=5\times 10^{-5}$ (with warm up and decay), WeightDecay=0.1, BatchSize=16, Epoch=5. The dataset is split into train-val with a 90-10 ratio based on the prompt level splitting and aiming for balanced scores in each split. We treat 0.5 as False and calculate negative scores as a binary metric. Results are presented in Table 7 with comparison with the same model without finetuning and LLaMA-Maverick.